Circular No. 7254

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GM SAGITTARII AND SAX J1819.3–2525 = XTE J1819–254 K. Ayani, Bisei Astronomical Observatory (BAO); and T. C. Peiris,

Arthur C. Clarke Institute for Modern Technologies, Sri Lanka, report: “A spectrogram (range 390–760 nm; resolution 0.7 nm at Hα) of GM Sgr was obtained with the BAO 1.01-m telescope on Sept. 15.50–15.52 UT, when the x-ray flux of SAX J1819.3–2525 was rapidly increasing (IAUC 7253). Broad emission lines of the Balmer series and of He II 468-nm are prominent. The Hα emission line has a FWHM of ∼ 2700 km/s and an equivalent width of 8 nm. The He II 541-nm emission line is also present. This suggests that GM Sgr is indeed the optical counterpart of SAX J1819.3–2525.”

R. M. Hjellming and M. P. Rupen, National Radio Astronomy Observatory; and A. J. Mioduszewski, Sydney University, report the Very Large Array detection on Sept. 16.02 UT of a radio source at α = 18^{h}19^{m}21^{s}.64, δ = −25^{o}24′25′′.6 (equinox 2000.0; uncertainty ± 0′′.3) with preliminary fluxes of 0.35, 0.4, and 0.25 Jy at 1.46, 4.9, and 14.9 GHz, respectively. This is 0′′.8 from the 1′′-uncertainty position of GM Sgr reported by Williams on IAUC 7253. There was no radio source above 5 mJy at this position in the 1.49-GHz NVSS image of the field in June 1996. This confirms the identification of SAX J1819.3–2525 = XTE J1819–254 (IAUC 7119, 7120, 7253) with the ‘cataclysmic variable’ GM Sgr.

M. Tichý and J. Tichá, Kleť Observatory, report that a CCD image taken on Sept. 15.839 UT, with cirrus clouds and GM Sgr at low altitude (limiting mag ∼ 13), shows GM Sgr at mag ∼ 11.5, indicating the star’s apparent rapid fading from the peak visual mag ∼ 9 noted on IAUC 7253.

W. Liller, Viña del Mar, Chile, reports that a low-dispersion spectrogram (∼ 0.7 nm/pixel at Hα) of GM Sgr taken with a CCD on Sept. 16.158 UT shows a strong Hα emission line at least 3.0 times the intensity of the neighboring continuum. (The spectrum is partially blended with that of nearby faint star.) Liller further reports that the broadband V magnitude on Sept. 16.131 was 13.27.

V1493 AQUILAE CCD magnitude estimates by D. Hanžl at Brno, Czech Republic (com-

parison star GSC 1048.863, assuming V = 8.40 ± 0.02, B − V = +0.311 ± 0.02, from Tycho Catalogue): Aug. 19.864 UT, V = 12.20 ± 0.05; 21.830, V = 11.51 ± 0.02; 22.840, V = 11.90 ± 0.01, B − V = +1.03 ± 0.01; Sept. 4.806, V = 13.92 ± 0.01, B − V = +0.80 ± 0.02; 9.788, V = 14.66 ± 0.05, B − V = +0.65 ± 0.05; 12.798, V = 14.82 ± 0.01, B − V = +0.64 ± 0.04.

1999 September 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green